OpenStudy (anonymous):

I need to find the slope (0,2) (-1,-4)

OpenStudy (saifoo.khan):

\[Slope = \frac{y_2-y_1}{x_2-x_1}\]

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (-4-2)/(-1-0) m = -6/(-1) m = 6